OpenStudy (anonymous):

Expand (5-5i)^4 Could someone walk me through this?

OpenStudy (anonymous):

U simply represent this like this from exonenets addition and subtraction rules : \[(5-5i)^2*(5-5i)^2\] Each is then exanded to : \[(5^2 - 50i + 25 i^{2} )\] which is simply : 25 - 50i - 25 = -50i then the original equation simplify to : -50i * -50i = \[250i^2 = -250\]

OpenStudy (anonymous):

Let me know if you got it :)

OpenStudy (anonymous):

I understand what you did but there is an answer key and it says the answer is \[2500(\cos \frac{ 28\Pi }{ 4 } + i \sin \frac{ 28\Pi }{ 4 }) =1.511-.450i\] how would i get to here

OpenStudy (amorfide):

you can use z=5-5i work this out in the form z=r(cosx+isinx) then z^4 would be z=r^4(cos4x+isin4x)

OpenStudy (amorfide):

\[z^{4}=r^{4}(\cos4x+isin4x)\]

OpenStudy (amorfide):

fixed it

OpenStudy (jdoe0001):

hmm

OpenStudy (amorfide):

So r is the magnitude of your imaginary number and x is the inverse tan of b/a where b is the coefficient of imaginary number and a is the real part

OpenStudy (amorfide):

\[r=\sqrt{(5^{2} + (-5)^{2}}\] to be clearer

OpenStudy (amorfide):